# On Three-Layer Data Markets

### v2

@@ 2.1 Utility functions
Measuring the revealed information, or information gain, through the variance of the posterior distribution has been used widely in economics, statistics, and learning theory (see, for example, Breiman et al. [1984], Bergemann et al. [2020], and Acemoglu et al. [2022]).
@@
is the user’s revealed information to platform $i$ if the user shares her information (i.e., when $a_{i}=1$) and the platform is offering service to the user (i.e., when $e_{i}=1$). The term
$\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},\bm{b})\triangleq\mathcal{I}(\bm{y}^{\top}\bm{\theta}~{}|~{}\bm{\tilde{s}}),$
is the user’s revealed information to the data buyer. Here, $\bm{\tilde{s}}$ denotes the vector $(\tilde{s}_{i})_{i=1}^{K}$, which is the revealed information to the data buyer. Notice that the user’s utility does not directly depend on $\bm{p}$, but we retain this notation to highlight that through the data buyer’s decision, the revealed information depends on $\bm{p}$. Also, notice that in finding the revealed information to the data buyer, all that matters is the set $\{\sigma_{i}~{}:~{}\text{ s.t. }a_{i}=b_{i}=e_{i}=1\}$. Therefore, we also find it convenient to define
@@
When the platform has not entered the market (i.e., $e_{i}=0$), the platform’s utility is zero. When the platform has entered the market (i.e., $e_{i}=1$), the second term $b_{i}p_{i}$ represents the platform’s revenue from selling the user’s data to the data buyer. The first term $a_{i}\mathcal{I}_{i}$, similar to the user’s utility, reflects the quality of service the platform provides (if the user ends up utilizing their service). This term is concluded to capture the fact that a better service by the platform also enhances the platform’s utility. For example, it could increase ad revenue or expand the user base. Lastly, the cost $c_{i}\in\mathbb{R}_{+}$ reflects the per-user cost incurred by the platform for providing the service.
@@ 2.2 Revealed Information Properties
Before analyzing this game, we provide some properties of our notion of revealed information. In stating the properties, we recall that the leaked information to the data buyer depends on $\bm{e}$, $\bm{a}$, and $\bm{b}$ through the set of platforms that share user data with the data buyer and the noise level of each of them. Therefore, we make use of the shorthand notation for the leaked information given in (2). In what follows, we also consider the lattice $(\mathcal{L},\leq)$, where
:={|$\mathcal{L}:=\{\ell~{}|~{}\ell\in\{0,1\}^{K}\}\text{ and }\ell\leq\ell^{\prime}\text{ if }\ell_{i}\leq\ell^{\prime}_{i}\text{ for all }i\in[K].$ (5)
@@
where $m_{i}=\gamma_{i}$, $[M]_{ii}=2+\sigma_{i}^{2}$, and $[M]_{ij}=\gamma_{i}\gamma_{j}$ for all $i,j\in S$ such that $i\neq j$.
Lemma 1 follows by using the properties of multivariate normal distribution. We should highlight that for simplicity, we assumed that the user adds the same noise level to their data before sharing it with any platform, which results in $\mathcal{I}_{i}$ for all $I\in[K]$ being equal. Our analysis continues to hold for any heterogeneous level of noise.
@@
The revealed information is increasing in the set of platforms that share the user data with the data buyer and is decreasing in $\bm{\sigma}$, i.e., for any $\bm{\sigma}$ and $S\subseteq[K]$, $\mathcal{I}(\bm{\sigma}_{S})$ is increasing in $S$ and decreasing in $\bm{\sigma}$.
@@
For a given vector of noise levels $\bm{\sigma}$, the revealed information is submodular in the set of platforms that share the user data with the data buyer, i.e., for any $\bm{\sigma}$, $S\subseteq T\subseteq[K]$, and $i\not\in T$, we have
@@
Lemma 4 states the data of platform $i$ reveals a relatively higher amount of information for smaller $\sigma_{i}$ values and for larger $\sigma_{j}$ values. To understand the former, let us compare the case of $\sigma_{i}\approx 0$ and $\sigma_{i}\to\infty$. When $\sigma_{i}\approx 0$, the data of platform $i$ reveals a lot, which helps the data buyer learn much better. However, when $\sigma_{i}\to\infty$, the data of platform $i$ does not contain any information, and therefore, the relative gain in the revealed information is $\approx 0$. To understand the former, let us again compare the case of $\sigma_{j}\approx 0$ and $\sigma_{j}\to\infty$. When $\sigma_{j}\approx 0$, the data of platform $j$ reveals a lot; therefore, the marginal increase in the revealed information of platform $i$ is small. However, when $\sigma_{j}\to\infty$, the data of platform $j$ does not contain any information, and therefore, the relative increase in the revealed information from platform $i$’s data is much larger.
@@ 3 Preliminary Equilibrium Characterization
$\mathcal{U}_{i}(\bm{\sigma},\bm{e},\bm{a},\bm{p}^{\mathrm{E}},\bm{b}^{\mathrm{E}})\geq\mathcal{U}_{i}(\bm{\sigma},\bm{e},\bm{a},(p_{i},\bm{p}^{\mathrm{E}}_{-i}),\bm{b})\text{ for all }i\in[K],\bm{b}^{\mathrm{E}}\in\mathcal{B}^{\mathrm{E}}(\bm{\sigma},\bm{e},\bm{a},\bm{p}^{\mathrm{E}}),p_{i},\bm{b}\in\mathcal{B}^{\mathrm{E}}(\bm{\sigma},\bm{e},\bm{a},(p_{i},\bm{p}^{\mathrm{E}}_{-i})).$
Before proceeding with the rest of the game, we establish that the data buyer and price equilibria admit a simple characterization, stated next.
@@
The proof of Proposition 1, given in the appendix, directly follows from the submodularity of the revealed information established in Lemma 3 for the following reason. Consider platform $i\in[K]$ and suppose the contrary that in equilibrium, the buyer is not purchasing from it. This platform has a profitable deviation by lowering its price so that the data buyer finds it optimal to purchase from it. Also, notice that the submodularity of the revealed information implies that if one more platform’s data is shared with the data buyer, the marginal increase in the revealed information decreases. Therefore, if the data buyer was purchasing from another platform $j\neq i$, they still find it optimal to purchase from that platform.
@@
$\Sigma_{\mathbf{\bm{a}},\bm{e}}:=\{(\sigma_{i}~{}:~{}i\text{ s.t.}e_{i}=1)~{}:~{}\bm{a}\text{ is the highest element of }\mathcal{A}^{\mathrm{E}}(\bm{\sigma},\bm{e})\}\text{ for all }\bm{a}\in\{0,1\}^{K}.$ (6)
@@ 4 Equilibrium Characterization with $K=2$ Platforms
[…]
@@
[…]
@@
Both parts of the result provide a lower bound on $\beta$, with the difference that the first part establishes a necessary condition for an equilibrium to exist, while the second part presents a sufficient condition. Note that the ratio between these two lower bounds converges to one as $\alpha$ grows. However, we next highlight a subtle difference between the two results. Note that, for any value $\xi$ greater than $2~{}(\max{c_{1},c_{2}}-\frac{1}{2})$, we can choose sufficiently large $\alpha$ and $\beta$ such that an equilibrium exists for these parameters and the ratio of $\beta$ over $\alpha$ is less than or equal to $\xi$. In other words, we have
@@
However, the infimum mentioned above does not imply that any $\alpha$ and $\beta$ satisfying the condition (11) will necessarily result in an equilibrium. In fact, as our proof establishes, to have an equilibrium for $\alpha=\bar{\alpha}$, we must have $\beta\geq\bar{\beta}(\bar{\alpha})$. In other words, we have
$\sup_{\alpha\geq\bar{\alpha}}\inf\left\{\frac{\beta}{\alpha}~{}\Big{|}~{}\beta\text{ s.t. an equilibrium exists with the given }\alpha\right\}=\bar{\beta}(\bar{\alpha}).$
@@
which implies $\sigma_{1}^{2}=2\alpha\gamma_{1}^{2}-2$. Notice that this is equivalent to the $\sigma_{1}$ corresponding to points $B$ and $X$ in Figure 1(b).
Now, for this noise level to be an equilibrium, we need to verify two potential deviations: (i) platform one has no incentive to abstain from participating, i.e., their utility should be nonnegative, and (ii) the second platform has no profitable deviation by entering the market. Using (13), the first condition simply implies
@@
Let us denote the noise levels corresponding to point $B$ as $(\sigma_{1}^{B},\sigma_{2}^{B})$. To understand why we only need to check the point $B$, consider that, given platform one’s chosen noise level (equal to $\sigma_{1}^{B}$), if the second platform enters the market and selects a certain noise variance, it will result in a point on the line $BX$. The smallest noise level for the second platform that motivates the user to share their data with them (and thus has the potential to be profitable) is $\sigma_{2}^{B}$. One might argue that the second platform could opt for a very large noise level so that the vector of noise levels falls into the green region $\tilde{\Sigma}_{(0,1)}$, implying that only the second platform receives the data. However, this is not feasible, as the line $BX$ stays within the blue region $\tilde{\Sigma}_{(1,1)}$ for any value of $\sigma_{2}$ (and in fact, it becomes tangent to the border of $\tilde{\Sigma}_{(0,1)}$ and $\tilde{\Sigma}_{(1,1)}$ as $\sigma_{2}$ approaches infinity). To confirm this, we need to demonstrate that
@@
Finally, we ask whether there is an equilibrium in which no platform enters the market. In such a scenario, we need to ensure no platform has a profitable deviation by entering the market and choosing some noise level. For instance, if platform $1$ wants to deviate, the smallest noise level that they would choose would be $\sigma_{1}^{B}$ as it makes the user indifferent between sharing and not sharing their data with them. In this case, the platform’s utility, as derived in (14), is given by
@@ 5 Equilibrium Characterization with $K>2$ Platforms
The comparison of the cost to $1/2$ in the above definition comes from the fact that the leaked information to any platform is $\mathcal{I}_{i}=1/2$ and therefore, high-cost platforms enter the market only if the payment they receive from the buyer is large enough. Conversely, low-cost platforms enter the market irrespective of the payment they receive from the buyer. We can view low-cost platforms as large platforms with small operation costs that do not need extra payments from the data buyer to enter the market and provide services. High-cost platforms, however, are small platforms with large operation costs that do need extra payments from the buyer to be able to provide service.
@@
Let us explain the qualifiers of Theorem 1 and then explain its intuition. First, notice that similar to the setting with $K=2$ platforms, for small enough $\alpha$ and also for intermediary values of $\beta$, a pure-strategy equilibrium may not exist. Therefore, we focus on large enough $\alpha$ and also avoid intermediary values of $\beta$. In the proof of Theorem 1, we have provided explicit expressions for $\bar{\alpha}$, $\bar{\beta}$, and $\underline{\beta}$.
To gain the intuition of the first part of Theorem 1, notice that for large enough $\beta$ the data buyer gains a lot when the user’s data is revealed to them and therefore is willing to pay a high price for the user data. This, in turn, means that even high-cost platforms find it optimal to enter the market. As established in Proposition 1 and Proposition 2, in equilibrium, the user shares with all platforms, and the data buyer purchases from all platforms. To see that the user utility becomes zero in equilibrium, suppose the contrary, that the user utility is strictly positive. This means that the equilibrium noise level strategy is in the interior of the set $\Sigma_{\mathbf{1},\mathbf{1}}$. Given that noise level strategy is in the interior of $\Sigma_{\mathbf{1},\mathbf{1}}$, there exists $i\in[K]$ such that by decreasing its noise level, we remain in the set $\Sigma_{\mathbf{1},\mathbf{1}}$ (i.e., the user still finds it optimal to share with all platforms), but this deviation increases the payment and therefore the utility of platform $i$. Also, Lemma 3 guarantees that the user still finds it optimal to share with all platforms.
@@
[…]
@@
6 Regulations
In this section, we consider regulations of the three-layer data market to improve the user utility. As we have established in Theorem 1, the user utility in equilibrium is zero. We now explore whether we can improve the user utility by imposing a regulation. As discussed in the introduction, one possible regulation is to impose a limit on the amount of information each platform can leak about the user to the data buyer. This means imposing a lower bound on the noise level. Let us formally define this regulation.

### v3

@@ 2.1 Utility functions
Measuring the revealed information, or information gain, through the variance of the posterior distribution has been used widely in economics, statistics, and learning theory [see, e.g., Breiman et al., 1984, Bergemann et al., 2020, Acemoglu et al., 2022].
@@
is the user’s revealed information to platform $i$ if the user shares her information (i.e., when $a_{i}=1$) and the platform is offering service to the user (i.e., when $e_{i}=1$). [^3] The user’s utility is additive across platforms, reflecting environments where the services are not substitutes and the user can engage with multiple platforms simultaneously. The term
[…]
is the user’s revealed information to the data buyer. [^4] The match between the user and the platform is captured by the term $x_{i}^{T}\theta$ and the user’s gain is the platform’s ability to learn (and act on) the match to deliver higher-quality service. Here, $\bm{\tilde{s}}$ denotes the vector $(\tilde{s}_{i})_{i=1}^{K}$, which is the revealed information to the data buyer. Notice that the user’s utility does not directly depend on $\bm{p}$, but we retain this notation to highlight that through the data buyer’s decision, the revealed information depends on $\bm{p}$. Also, notice that in finding the revealed information to the data buyer, all that matters is the set $\{\sigma_{i}~:~\text{ s.t. }a_{i}=b_{i}=e_{i}=1\}$. Therefore, we also find it convenient to define
@@
When the platform has not entered the market (i.e., $e_{i}=0$), the platform’s utility is zero. When the platform has entered the market (i.e., $e_{i}=1$), the second term $b_{i}p_{i}$ represents the platform’s revenue from selling the user’s data to the data buyer. The first term $a_{i}\mathcal{I}_{i}$, similar to the user’s utility, reflects the quality of service the platform provides (if the user ends up utilizing their service). This term is included to capture the fact that a better service by the platform also enhances the platform’s utility. For example, it could increase ad revenue or expand the user base. Lastly, the cost $c_{i}\in\mathbb{R}_{+}$ reflects the per-user cost incurred by the platform for providing the service.
@@
Remark 1.
In our baseline model, we assume the utility of all players to be linear in the leaked information. In Appendix B.1, we establish that our main results continue to hold more broadly. Specifically, we let the user utility be [^5] We retain a linear form for the first $\mathcal{I}_{i}$ because this is a constant term and applying a function to it does not change the analysis.
$\mathcal{U}_{\text{user}}(\bm{\sigma},\bm{e},\bm{a},\bm{p},\bm{b}):=\sum_{i=1}^{K}a_{i}e_{i}\mathcal{I}_{i}-\alpha~H_{u}\left(\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},\bm{b})\right),$ (5)
and the data buyer’s utility be
$\mathcal{U}_{\text{buyer}}(\bm{\sigma},\bm{e},\bm{a},\bm{p},\bm{b}):=\beta~H_{b}\left(\mathcal{I}(\bm{\sigma},\bm{e},\bm{a},\bm{b})\right)-\sum_{i=1}^{K}b_{i}p_{i}.$ (6)
We then prove that our results continue to hold if the functions $H_{u}(\cdot)$ and $H_{b}(\cdot)$ belong to the following family of functions:
$\mathcal{H}=\{H:[0,1]\mapsto[0,1]~:~\inf_{x\in[0,1]}H^{\prime}(x)>0,~H(\cdot)\text{ is concave},~\text{ and }H(0)=0\}.$
@@ 2.2 Revealed Information Properties
Before analyzing this game, we state some properties of our notion of revealed information. We recall that the leaked information to the data buyer depends on $\bm{e}$, $\bm{a}$, and $\bm{b}$ through the set of platforms that share user data with the data buyer and the noise level of each of them. Therefore, we make use of the shorthand notation for the leaked information given in (2). In what follows, we also consider the lattice $(\mathcal{L},\leq)$, where
$\mathcal{L}:=\{\ell~|~\ell\in\{0,1\}^{K}\}\text{ and }\ell\leq\ell^{\prime}\text{ if }\ell_{i}\leq\ell^{\prime}_{i}\text{ for all }i\in[K].$ (7)
@@
where $m_{i}=\gamma_{i}$, $M]_{ii}=2+\sigma_{i}^{2}$, and $M]_{ij}=\gamma_{i}\gamma_{j}$ for all $i,j\in S$ such that $i\neq j$.
Lemma 1 follows by using the properties of the multivariate normal distribution. [^6] For simplicity, we assumed that the user adds the same noise level with variance one to their data before sharing it with any platform, resulting in $\mathcal{I}_{i}=1/2$ for all $i\in[K]$. Our analysis continues to hold for any heterogeneous noise levels.
@@
The revealed information is increasing in the set of platforms that share the user data with the data buyer and is decreasing in $\bm{\sigma}$; i.e., for any $\bm{\sigma}$ and $S\subseteq[K]$, $\mathcal{I}(\bm{\sigma}_{S})$ is increasing in $S$ and decreasing in $\bm{\sigma}$.
@@
For a given vector of noise levels $\bm{\sigma}$, the revealed information is submodular in the set of platforms that share the user data with the data buyer; i.e., for any $\bm{\sigma}$, $S\subseteq T\subseteq[K]$, and $i\not\in T$, we have
@@
Lemma 4 states that the data of platform $i$ reveals a relatively larger amount of information for smaller values of $\sigma_{i}$ and for larger values of $\sigma_{j}$. To understand the former, let us compare the case of $\sigma_{i}\approx 0$ and $\sigma_{i}\to\infty$. When $\sigma_{i}\approx 0$, the data of platform $i$ reveals a lot, which helps the data buyer learn better. However, when $\sigma_{i}\to\infty$, the data of platform $i$ does not contain any information, and therefore, the relative gain in the revealed information is $\approx 0$. To understand the latter, let us again compare the case of $\sigma_{j}\approx 0$ and $\sigma_{j}\to\infty$. When $\sigma_{j}\approx 0$, the data of platform $j$ reveals a lot; therefore, the marginal increase in the revealed information of platform $i$ is small. However, when $\sigma_{j}\to\infty$, the data of platform $j$ does not contain any information, and therefore, the relative increase in the revealed information from platform $i$’s data is much larger.
The lemmas above, which establish various properties of the revealed information, allow us to characterize the equilibrium properties in the next section. Although the derivation of these lemmas relies on Equation 35 and the definition of revealed information, one can relax that assumption or adopt an alternative notion of information gain, and our results on the equilibrium properties would continue to hold as long as the aforementioned properties remain valid.
@@ 3 Preliminary Equilibrium Characterization
$\mathcal{U}_{i}(\bm{\sigma},\bm{e},\bm{a},\bm{p}^{\mathrm{E}},\bm{b}^{\mathrm{E}})\geq\mathcal{U}_{i}(\bm{\sigma},\bm{e},\bm{a},(p_{i},\bm{p}^{\mathrm{E}}_{-i}),\bm{b})$
$\text{ for all }i\in[K],\bm{b}^{\mathrm{E}}\in\mathcal{B}^{\mathrm{E}}(\bm{\sigma},\bm{e},\bm{a},\bm{p}^{\mathrm{E}}),p_{i},\bm{b}\in\mathcal{B}^{\mathrm{E}}(\bm{\sigma},\bm{e},\bm{a},(p_{i},\bm{p}^{\mathrm{E}}_{-i})).$
Before proceeding with the rest of the game, we establish that the data buyer and price equilibria admit a simple characterization, as stated next.
@@
To see why this result holds, first note that for any platform $i$ that has entered the market and with which the user has shared data (i.e., $e_{i}=a_{i}=1$), we have $b_{i}=1$ in equilibrium. This is because if that were not the case, the platform would have a profitable deviation by setting a very low but positive price for its data to incentivize the buyer to purchase it. This positive price exists because, if the buyer obtains the platform’s data, their utility increases; thus, any price below that positive utility gain multiplied by $\beta$ would still incentivize the buyer to buy the data.
Next, knowing that all these $b_{i}$’s are equal to one in equilibrium, we claim that the prices given in the statement of Proposition 1 constitute an equilibrium. The proof is straightforward: no platform has an incentive to decrease this price (since the buyer would still be willing to pay more), nor to increase it (since the buyer would no longer purchase from them). The proof of uniqueness is provided in the Appendix.
@@
[…]
@@ 4 Equilibrium Characterization with $K=2$ Platforms
Figure 3: Four possible sets $\tilde{\Sigma}_{\bm{a}}$ for $\bm{a}\in\{0,1\}^{2}$ for a setting with $K=2$ platforms, $\gamma_{1}=.8$, and $\gamma_{2}=.7$.
@@
$\alpha\geq\bar{\alpha}\quad\text{and}\quad\beta\geq 2\alpha~(\max\{c_{1},c_{2}\}-\frac{1}{2}).$ (13)
@@
Both parts of the result provide a lower bound on $\beta$, with the difference that the first part establishes a necessary condition for an equilibrium to exist, while the second part presents a sufficient condition. Note that the ratio between these two lower bounds converges to one as $\alpha$ grows. However, we next highlight a subtle difference between the two results. Note that, for any value $\xi$ greater than $2~(\max{c_{1},c_{2}}-\frac{1}{2})$, we can choose $\alpha$ and $\beta$ sufficiently large such that an equilibrium exists for these parameters and the ratio of $\beta$ over $\alpha$ is less than or equal to $\xi$. In other words, we have
@@
However, the form of the infimum above does not imply that any $\alpha$ and $\beta$ satisfying the condition (13) will necessarily result in an equilibrium. In fact, as our proof establishes, to have an equilibrium for $\alpha=\bar{\alpha}$, we must have $\beta\geq\bar{\beta}(\bar{\alpha})$. In other words, we have
[…]
@@
which implies $\sigma_{1}^{2}=2\alpha\gamma_{1}^{2}-2$. Notice that this is equivalent to the values of $\sigma_{1}$ corresponding to points $B$ and $X$ in Figure 3(b).
Now, for this noise level to be an equilibrium, we need to verify two potential deviations: (i) platform 1 has no incentive to abstain from participating, i.e., their utility should be nonnegative, and (ii) platform 2 has no profitable deviation by entering the market. Using (15), the first condition simply implies
@@
Let us denote the noise levels corresponding to point $B$ as $(\sigma_{1}^{B},\sigma_{2}^{B})$. To understand why we only need to check the point $B$, consider that, given platform 1’s chosen noise level (equal to $\sigma_{1}^{B}$), if the second platform enters the market and selects a certain noise variance, it will result in a point on the line $BX$. The smallest noise level for the second platform that motivates the user to share their data with them (and thus has the potential to be profitable) is $\sigma_{2}^{B}$. One might argue that the second platform could opt for a very large noise level so that the vector of noise levels falls into the green region $\tilde{\Sigma}_{(0,1)}$, implying that only the second platform receives the data. However, this is not feasible, as the line $BX$ stays within the blue region $\tilde{\Sigma}_{(1,1)}$ for any value of $\sigma_{2}$ (and in fact, it becomes tangent to the border of $\tilde{\Sigma}_{(0,1)}$ and $\tilde{\Sigma}_{(1,1)}$ as $\sigma_{2}$ approaches infinity). To confirm this, we need to demonstrate that
@@
Finally, we ask whether there is an equilibrium in which no platform enters the market. In such a scenario, we need to ensure that no platform has a profitable deviation by entering the market and choosing some noise level. For instance, if platform $1$ wants to deviate, the smallest noise level that they would choose would be $\sigma_{1}^{B}$ as it makes the user indifferent between sharing and not sharing their data with them. In this case, the platform’s utility, as derived in (16), is given by
@@ 5 Equilibrium Characterization with $K>2$ Platforms
The comparison of the cost to $1/2$ in the above definition comes from the fact that the leaked information to any platform is $\mathcal{I}_{i}=1/2$ and therefore, high-cost platforms enter the market only if the payment they receive from the buyer is large enough. Conversely, low-cost platforms enter the market irrespective of the payment they receive from the buyer. We can view low-cost platforms as those that do not necessarily need additional payments from the data buyer to enter the market and provide services. These platforms might include large ones with small operating costs or those that do not rely solely on selling data to generate profit. On the other hand, high-cost platforms are the opposite; they would require some payment from the data buyer to be able to provide services.
@@
Let us explain the qualifiers of Theorem 1 and then explain its intuition. First, notice that, similar to the setting with $K=2$ platforms, for small enough $\alpha$ and also for intermediary values of $\beta$, a pure-strategy equilibrium may not exist. Therefore, we focus on large enough $\alpha$ and also avoid intermediary values of $\beta$. In the proof of Theorem 1, we have provided explicit expressions for $\bar{\alpha}$, $\bar{\beta}$, and $\underline{\beta}$. In particular, $\bar{\alpha}=(K+1)^{2}/8$ depends only on $K$ while $\underline{\beta}$ and $\bar{\beta}$ depend on the model parameters $\{(c_{i},\gamma_{i})\}_{i\in[K]}$.
To gain the intuition of the first part of Theorem 1, notice that for large enough $\beta$ the data buyer gains a lot when the user’s data is revealed to them and therefore is willing to pay a high price for the user data. This, in turn, means that even high-cost platforms find it optimal to enter the market. As established in Proposition 1 and Proposition 2, in equilibrium, the user shares with all platforms, and the data buyer purchases from all platforms. To see that the user utility becomes zero in equilibrium, suppose the contrary, that the user utility is strictly positive. This means that the equilibrium noise level strategy is in the interior of the set $\Sigma_{\mathbf{1},\mathbf{1}}$. Given that the noise level strategy is in the interior of $\Sigma_{\mathbf{1},\mathbf{1}}$, there exists $i\in[K]$ such that by decreasing its noise level, we remain in the set $\Sigma_{\mathbf{1},\mathbf{1}}$ (i.e., the user still finds it optimal to share with all platforms), but this deviation increases the payment and therefore the utility of platform $i$. Also, Lemma 3 guarantees that the user still finds it optimal to share with all platforms.
@@
Remark 2.
While Theorem 1 focuses on the equilibrium characterization for small and sufficiently large values of $\beta$, in Section B.2 we further characterize the equilibrium for intermediate values of $\beta$. In particular, we establish the existence of a sequence of equilibria in which the low-cost platforms always enter the market, and as $\beta$ increases, the high-cost platforms enter sequentially, in the order of their costs (see Theorem 2). Furthermore, we show that for each high-cost platform, there exists a minimum value of $\beta$ required for market entry, and this threshold increases linearly with the platform’s cost of production (see Proposition 9).
@@
    $(\text{number of low-cost platforms})\left(\frac{1}{2}+\frac{\beta}{2\alpha}\right)-\sum_{i\in[K]~:~c_{i}<\frac{1}{2}}c_{i}.$
@@
6 Regulation
In this section, we consider regulation of the three-layer data market to improve the user utility. As we have established in Theorem 1, the user utility in equilibrium is zero. We now explore whether we can improve the user utility by imposing some form of regulation. As discussed in the introduction, one possible regulation is to impose a limit on the amount of information each platform can leak about the user to the data buyer. This means imposing a lower bound on the noise level. Let us formally define this regulation.